OpenStudy (anonymous):

Rationalize the denominator of 3 over√2??

OpenStudy (across):

\[\frac{3}{\sqrt{2}}=\frac{3}{\sqrt{2}}\cdot\frac{\sqrt{2}}{\sqrt{2}}=\frac{3\sqrt{2}}{2}\]